# Galaxy And Mass Assembly (GAMA): from filaments to voids, how extreme environment affects gas metallicity and SFR in galaxies

### v1

@@ Abstract
Context. The stellar mass-metallicity (–) and stellar mass-star formation rate (–SFR) relations are fundamental tools for understanding the evolution of star-forming (SF) galaxies. Combined with environmental factors, these relations provide valuable insights into how galaxies evolve.
Aims. We analyse the – and –SFR relations for SF galaxies, classified according to their environment, and to compare them with a control sample of field galaxies. The aim is to quantify the differences in metallicity () and star formation rate (SFR) among galaxies in different environments. To achieve this, we are using data from the Galaxy And Mass Assembly (GAMA) survey, along with the filament catalogue that classifies galaxies into filaments, tendrils, and voids.
@@
Results. We determine that increases as environments become denser. On the contrary, SFR increased as environments became less dense. Both results demonstrate significant differences between filaments and tendrils compared to voids. We also analyse galaxies in filaments and tendrils that do not belong to any group, finding little to no difference compared to their control sample. Morphology showed no significant deviation from the control sample.
Conclusions. We find that galaxies in filaments have higher metallicities and lower SFRs due to enriched environments, while void galaxies sustain high SFRs with low metallicities, likely driven by isolation and cold gas accretion. Our results indicate that local environmental factors, rather than large-scale structure, are the primary drivers of these differences.
@@
Over the past decades, various empirical methods have been developed to estimate gas metallicities (Z) from galaxy emission lines, using different elements as tracers (e.g., Poetrodjojo et al. 2019). Moreover, metallicity has been found to correlate with stellar mass, giving rise to the well-known – relation. The study of this correlation has evolved and adapted to different galaxy types, providing valuable insights into its behaviour (Lequeux et al. 1979; Tremonti et al. 2004; Sánchez-Menguiano et al. 2024a, 2024b; Wu & Zhang 2024).
@@
Similarly, the star formation rate (SFR) measures the amount of stellar mass formed per unit time, and several empirical approaches exist for its estimation. One of the most widely used is based on the H emission line. These methods have established a second key correlation, between SFR and stellar mass (–SFR relation), which has proven robust across various contexts (Guzmán et al. 1997; Buat et al. 2008; Davies et al. 2016; Thorne et al. 2021). Together, these relations reveal that more massive galaxies generally exhibit higher and SFR values, while lower-mass systems show lower values for both parameters. Furthermore, studies have proposed a combined relation involving stellar mass, metallicity, and SFR, providing a more comprehensive view of galaxy evolution (Lara-López et al. 2010, e.g.,, 2013).
@@
These correlations, however, have also been explored in broader contexts. One key aspect is morphology: the SFR appears to be influenced by galaxy structure (Porter-Temple et al. 2022), with spiral disks following distinct trends (Willett et al. 2015), while elliptical galaxies tend to exhibit higher metallicities (Pipino & Matteucci 2006; Pipino et al. 2011). Morphology itself has been linked to the galactic environment, with evidence suggesting that morphological type plays a more significant role than environment in determining parameter evolution (Calvi et al. 2018, e.g.,; Laishram et al. 2024). Nevertheless, environment remains a crucial factor, as shown by Sureshkumar et al. 2024, who found that galaxy mergers are more frequent in under-dense regions, impacting various galactic properties. Finally, the temporal evolution of these parameters has also been widely investigated (Noeske et al. 2007, e.g.,; Henry et al. 2021), particularly in the era of next-generation telescopes like JWST, which has revealed that at high redshifts, galaxy evolution remains strongly environment-dependent (Curti et al. 2024, e.g.,; Laishram et al. 2024). Lastly, it is important to highlight the current role of machine learning in the study of these parameters. Some machine learning studies are capable of deriving and SFR from galaxy images (Wu & Boada 2019, e.g.,; Alfonzo et al. 2024).
@@
Building on studies of environmental dependence in and SFR, several authors have further examined these relationships. Sotillo-Ramos et al. (2021) found that local density, rather than group-centric distance, more strongly affects both SFR and , indicating that stellar mass is likely the main factor driving quenching in star-forming galaxies. Similarly, Wijesinghe et al. (2012) argued that stellar mass primarily influences SFRs, with minimal environmental effects in star-forming galaxies. They also suggested that SFR-density trends are mainly driven by changes in galaxy morphology. On the other hand, Ellison et al. (2008) showed that for galaxies in pairs, the bulge fraction significantly affects . Finally, the quenching process described by Sotillo-Ramos et al. (2021), where red galaxies are often found in denser environments, has been linked to both and SFR in other studies, such as McNaught-Roberts et al. (2014). However, some reports suggest that processes like starvation or ram-pressure stripping may play a more prominent role than quenching (e.g., Trussler et al. 2020).
@@
In addition to these parameters, large-scale structure has also been studied. Many papers, such as those reported by de Lapparent et al. (1986) and Alpaslan et al. (2014b), have paved the way for classifying the distribution of galaxies in the Universe. Galaxies tend to cluster into groups and clusters ( 1 Mpc) that gravitationally clump together to form super-clusters ( 10 Mpc). These super-clusters appear to be linked together in the form of chains of galaxies, creating filaments ( 100 Mpc) that surround voids, which correspond to low-density regions of the Universe. Attached to these filaments are branches that protrude from these structures, known as ’tendrils’. These regions were first introduced by Alpaslan et al. (2014b). Based on these previous reports, galaxies can be classified into the following groups: galaxies in filaments, tendrils galaxies, and void galaxies.
@@
Regarding galaxies in filaments and field galaxies, O’Kane et al. (2024) concludes that the differences in star formation activity and morphology between galaxies in filaments and those in the field are primarily driven by local density, rather than the filaments themselves. Focusing on galaxies in voids, Porter et al. (2023) and Argudo-Fernández et al. (2024) argue that these galaxies are predominantly late-type, with a morphology that is independent of both void size and galaxy density. In addition to filaments and voids, Domínguez-Gómez et al. (2023a) also consider walls (dense regions of galaxies and dark matter that form flat or filament-like structures, connecting galaxy clusters and filaments). They all show that void galaxies tend to have slightly lower stellar metallicities compared to those in filaments and walls, and significantly lower than cluster galaxies. Besides, these differences are most evident in low-mass, extended-SFH, spiral, and blue galaxies. Despite these findings, many authors (Einasto et al. 2022; Hoosain et al. 2024; O’Kane et al. 2024; Torres-Ríos et al. 2024) emphasise that galaxy evolution is primarily driven by both local and large-scale environments. High-density regions lead to quenched star formation and early-type galaxies, while low-density regions show more varied star formation histories. Although some authors give greater weight to the local environment (Kauffmann et al. 2004, Vulcani et al. 2012, Robotham et al. 2013).
@@
The goal of this work is to examine the relation between star formation and the metallicity of galaxies and their position in this large-scale structure. We will use a well-determined sample with spectroscopic redshifts and ample ancillary data.
@@
This paper is structured as follows: Section 2 explains the Galaxy And Mass Assembly (GAMA) survey and how the filtering process is carried out to select the sample of galaxies that are going to be studied here. The fluxes of these filtered galaxies are corrected for dust extinction, and star-forming (SF) galaxies are selected using the Baldwin-Phillips-Terlevich (BPT) diagram. Metallicity and SFR are then derived from their emission lines. After, the different types of galaxies (filaments, tendrils, and voids) are explained, and control samples are generated from field galaxies of GAMA under specific established conditions and tests. In Section 3, the scaling relations found from the different samples are analysed and compared. For that, a classical and Bayesian analysis has been carried out to compare the differences and SFR of the galaxies in the different environments with a control sample. Finally, Section 4 discusses the results obtained, extracting the conclusions of this work in Section 5.
@@ 2.1 GAMA survey
The Galaxy and Mass Assembly (GAMA) (Driver et al. 2011; Liske et al. 2015) is a large survey whose primary objective is to map the distribution of galaxies at various scales, with a focus on obtaining high-quality spectroscopic data for each galaxy. The data were taken with the 3.9 m Anglo-Australian Telescope (AAT) employing the AAOmega multi-object spectrograph (Sharp et al. 2006) and the 2dF fibre feed. The spectra were taken using a 2 arcsec diameter fibre, covering a range between 3740Å to 8850Å, with a resolution of 3.2Å. The survey covers a large area of the sky, approximately 286 square degrees, and includes over 340 000 galaxies. This study will focus on the GAMA Phase 1, which collects 192 545 galaxies in 48 square degrees, and a limiting Petrosian magnitude of m19.8 mag in one field and m19.4 mag in the other two. For further details see Baldry et al. (2010), Driver et al. (2011), Hopkins et al. (2013), Liske et al. (2015).
On the other hand, the emission lines from GAMA were measured using the Gas AND Absorption Line Fitting (GANDALF) algorithm (Sarzi et al. 2006), which is a simultaneous emission and absorption line fitting algorithm. GANDALF measures and corrects for stellar absorption of the emission lines. All these line measurements include the emission lines flux, the equivalent width (EW), and the signal-to-noise ratio (S/N) for the most prominent emission lines, among other parameters. For more details, see Gordon et al. (2018).
The full GAMA catalogue is matched with another catalogue of stellar masses (Taylor et al. 2011), and filaments (Alpaslan et al. 2014a), both catalogues from the GAMA survey. The first one (Taylor et al. 2011) estimates the stellar mass-to-light ratio (/L) from optical photometry using stellar population synthesis based on the Bruzual & Charlot (2003) model. The second one (Alpaslan et al. 2014a) used an adapted minimal spanning tree algorithm called Friend of Friend (FoF) (Merson et al. 2013), which groups galaxies based on their proximity to one another, thus creating a network of connected galaxies. It identifies galaxies within a specified linking length and classifies them based on their connection to nearby galaxies.
@@
S/N 3 for H and [O iii] 5007 and S/N 2 for H and [N ii] 6583
0.3: to ensure the H emission line remains within the optical range, as it shifts out at higher redshifts
@@
Condition | Number of galaxies |
Initial sample | 192 545 |
After S/N cut (i) | 18 247 |
After redshift cut (ii) | 17 283 |
Therefore, the initial number of galaxies available in the GAMA catalogue for this study has been reduced from 192 545 to 17 283 after the filtering process.
Next, we corrected all the emission lines for dust extinction. The extinction curve used in this work is from Cardelli et al. (1989), we assume a case B of recombination (Osterbrock 1989) and = 2.86.
Therefore, the corrected flux is given by , where and correspond to corrected and observed flux, respectively. The parameter is the extinction for a given wavelength. When is lower than the theoretical value, then the adopted correction is set to zero.
@@ 2.2 BPT Diagram
The Baldwin-Phillips-Terlevich (BPT) diagram (Baldwin et al. 1981; Veilleux & Osterbrock 1987) is commonly used to classify and understand the ionisation sources of emission-line galaxies. This diagram will allow us to classify the galaxies from the GAMA survey into three groups: star-forming (SF), active galactic nucleus (AGN), and composite galaxies.
The BPT diagram of Figure 1 has been plotted using the emission lines of [O iii] 5007, [N ii] 6583, H and H from the 17 283 galaxies selected. We use the curve of Kauffmann et al. (2003) to separate composite from AGN galaxies, while the curve of Kewley et al. (2001) separates composite from SF galaxies.
The estimation of the error bars, shown in Figure 1, has been done using Monte Carlo propagation. This method generates simulated data for each variable from a multivariate t-distribution, using the means and the full covariance matrix, which includes both the diagonal elements (variances) and the off-diagonal terms (covariances) derived from the uncertainties of the original variables.
@@
Galaxy type | Number of galaxies |
SF | 13 689 |
Composite | 2 336 |
AGN | 1 258 |
The data of Table 2 indicates that a total of 79% of the sample (13 689 galaxies) consists of SF galaxies, which will be the focus of our analysis hereafter.
@@ 2.3 Metallicity and star formation rate estimation
Gas metallicities are estimated using the extinction-corrected fluxes of the star-forming galaxies, and the calibration method from Pettini & Pagel (2004).
@@
and is in the range of -1 1.9. We adopt the O3N2 calibration, as it provides reliable metallicity estimates for large samples of star-forming galaxies (Kewley & Ellison 2008). Additionally, since the emission-line ratios involved are close in wavelength, this method is minimally affected by dust extinction (Lara-López et al. 2013).
@@
where the luminosity of the H line, corrected for dust obscuration, is defined by Hopkins et al. (2003) as follows
where and are the observed fluxes of the H and H lines, respectively, corrected by stellar absorption by GANDALF. The EWHα denotes the equivalent width of the H line in angstroms. Finally, denotes the absolute Petrosian magnitude in the r-band. With these considerations, the representation of the stellar mass-metallicity (–) and stellar mass-star formation rate (–SFR) relation can be established.
It is important to notice that the fibre size of the AAT could introduce some bias in the estimation of and SFR parameters. This issue is particularly relevant for SFR measurements due to the galaxy’s angular size. If a galaxy is large, its outer regions might not fit entirely within the fibre, potentially leading to incomplete SFR measurements. However, this potential bias is mitigated by the calibration method used by Hopkins et al. (2003), which accounts for such discrepancies in Equation 4. In contrast, the parameter is less significantly affected. According to Kewley et al. (2005), as long as at least 20% of the light of the galaxy falls within the optical fibre, the gas metallicity provides a reasonable approximation of the total value.

### v2

@@ Abstract
Context. The stellar massmetallicity (–) and stellar massstar formation rate (–SFR) relations are fundamental tools for understanding the evolution of star-forming (SF) galaxies. Combined with environmental factors, these relations provide valuable insights into how galaxies evolve.
Aims. We analysed the – and –SFR relations for SF galaxies, classified according to their environment, and compared them with a control sample of field galaxies. The aim was to quantify the differences in metallicity () and star formation rate (SFR) among galaxies in different environments. To achieve this, we used data from the Galaxy And Mass Assembly (GAMA) survey, along with the filament catalogue that classifies galaxies into filaments, tendrils, and voids.
@@
Results. We determined that increases as environments become denser. On the contrary, SFR increases as environments become less dense. Both results demonstrate significant differences between filaments and tendrils compared to voids. We also analysed galaxies in filaments and tendrils that do not belong to any group, and found little to no difference compared to their control sample. Morphology showed no significant deviation from the control sample.
Conclusions. We find that galaxies in filaments and tendrils have higher metallicities and lower SFRs due to enriched environments, while void galaxies sustain high SFRs with low metallicities, likely driven by isolation and cold gas accretion. Our results indicate that local environmental factors, rather than large-scale structure, are the primary drivers of these differences.
@@
Over the past decades, various empirical methods have been developed to estimate gas metallicities (Z) from galaxy emission lines, using different elements as tracers (e.g. Poetrodjojo et al. 2019). Moreover, metallicity has been found to correlate with stellar mass, giving rise to the well-known massmetallicity (–) relation. The study of this correlation has evolved and adapted to different galaxy types, providing valuable insights into its behaviour (Lequeux et al. 1979; Tremonti et al. 2004; Sánchez-Menguiano et al. 2024a, 2024b; Wu & Zhang 2024).
@@
Similarly, the star formation rate (SFR) measures the amount of stellar mass formed per unit time, and several empirical approaches exist for its estimation. One of the most widely used is based on the H emission line. These methods have established a second key correlation, the massstar formation rate (–SFR) relation, which has proven robust across various contexts (Guzmán et al. 1997; Buat et al. 2008; Davies et al. 2016; Thorne et al. 2021). Together, these relations reveal that more massive galaxies generally exhibit higher and SFR values, while lower-mass systems show lower values for both parameters. Furthermore, studies have proposed a combined relation involving stellar mass, metallicity, and SFR, in order to provide a more comprehensive view of galaxy evolution (Lara-López et al. 2010, e.g., 2013).
@@
These correlations, however, have also been explored in broader contexts. One key aspect is morphology: the SFR appears to be influenced by galaxy structure (Porter-Temple et al. 2022), with spiral disks following distinct trends (Willett et al. 2015), while elliptical galaxies tend to exhibit higher metallicities (Pipino & Matteucci 2006; Pipino et al. 2011). Morphology itself has been linked to the galactic environment, with evidence suggesting that morphological type plays a more significant role than environment in determining parameter evolution (Calvi et al. 2018, e.g.; Laishram et al. 2024). Nevertheless, environment remains a crucial factor, as shown by Sureshkumar et al. 2024, who found that galaxy mergers are more frequent in underdense regions, and impact various galactic properties. Finally, the temporal evolution of these parameters has also been widely investigated (Noeske et al. 2007, e.g.; Henry et al. 2021), particularly in the era of next-generation telescopes like JWST, which has revealed that at high redshifts galaxy evolution remains strongly environment-dependent (Curti et al. 2024, e.g.; Laishram et al. 2024). Lastly, it is important to highlight the current role of machine learning in the study of these parameters. Some machine learning studies are capable of deriving and SFR from galaxy images (Wu & Boada 2019, e.g.; Alfonzo et al. 2024).
@@
Building on studies of environmental dependence in and SFR, several authors have examined these relationships further. Sotillo-Ramos et al. (2021) found that local density, rather than group-centric distance, more strongly affects both SFR and , indicating that stellar mass is likely the main factor driving quenching in star-forming galaxies. Similarly, Wijesinghe et al. (2012) argued that stellar mass primarily influences SFRs, with minimal environmental effects in star-forming galaxies. They also suggested that SFR-density trends are mainly driven by changes in galaxy morphology. On the other hand, Ellison et al. (2008) showed that for galaxies in pairs, the bulge fraction significantly affects . Finally, the quenching process described by Sotillo-Ramos et al. (2021), where red galaxies are often found in denser environments, has been linked to both and SFR in other studies, such as McNaught-Roberts et al. (2014). However, some reports suggest that processes such as starvation or ram-pressure stripping may play a more prominent role than quenching (e.g. Trussler et al. 2020), with the latter being one of the main mechanisms responsible for suppressing the inflow of metal-poor gas, thereby preventing dilution (Khoram et al. 2025).
@@
In addition to these parameters, large-scale structure has also been studied. Many papers, such as those reported by de Lapparent et al. (1986) and Alpaslan et al. (2014b), have paved the way for classifying the distribution of galaxies in the Universe. Galaxies tend to cluster into groups and clusters ( 1 Mpc) that gravitationally clump together to form super-clusters ( 10 Mpc). These super-clusters appear to be linked together in the form of chains of galaxies, creating filaments ( 100 Mpc) that surround voids, which correspond to low-density regions of the Universe. Attached to these filaments are branches that protrude from these structures, known as ‘tendrils’. These regions were first introduced by Alpaslan et al. (2014b). Based on these previous reports, galaxies can be classified into the following groups: galaxies in filaments, tendrils galaxies, and void galaxies.
@@
O’Kane et al. (2024) concludes that the differences in star formation activity and morphology between galaxies in filaments and those in the field are primarily driven by local density, rather than the filaments themselves. Focusing on galaxies in voids, Porter et al. (2023) and Argudo-Fernández et al. (2024) argue that these galaxies are predominantly late type, with a morphology that is independent of both void size and galaxy density. In addition to filaments and voids, Dom´ınguez-Gómez et al. (2023a) also consider walls (dense regions of galaxies and dark matter that form flat or filament-like structures, connecting galaxy clusters and filaments). They all show that void galaxies tend to have slightly lower stellar metallicities compared to those in filaments and walls, and significantly lower than cluster galaxies. In addition, these differences are most evident in low-mass, extended-SFH, spiral, and blue galaxies. Despite these findings, many authors (Einasto et al. 2022; Hoosain et al. 2024; O’Kane et al. 2024; Torres-R´ıos et al. 2024) emphasise that both local and large-scale environments primarily drive galaxy evolution. High-density regions lead to quenched star formation and early-type galaxies, while low-density regions show more varied star formation histories. Although some authors give greater weight to the local environment (Kauffmann et al. 2004; Vulcani et al. 2012; Robotham et al. 2013).
@@
The goal of this work is to examine the relation between star formation and the metallicity of galaxies and their position in this large-scale structure. To this end, we used a well-determined sample with spectroscopic redshifts and ample ancillary data.
@@
This paper is structured as follows. Section 2 explains the Galaxy And Mass Assembly (GAMA) survey and how the filtering process was carried out to select the sample of galaxies studied here. The fluxes of these filtered galaxies were corrected for dust extinction, and star-forming (SF) galaxies were selected using the Baldwin-Phillips-Terlevich (BPT) diagram. Metallicity and SFR were then derived from their emission lines. Afterwards, the different types of galaxies (filaments, tendrils, and voids) are explained, and control samples are generated from field galaxies of GAMA under specific established conditions and tests. In Section 3, the scaling relations found from the different samples are analysed and compared. For this purpose, a classical and Bayesian analysis was conducted to compare the differences and SFR of the galaxies in the various environments with a control sample. Finally, Section 4 discusses the results obtained, and we present the conclusions of this work in Section 5.
@@ 2.1 GAMA survey
The Galaxy and Mass Assembly (GAMA) (Driver et al. 2011; Liske et al. 2015) is a large survey whose primary objective is to map the distribution of galaxies at various scales, with a focus on obtaining high-quality spectroscopic data for each galaxy. The data were taken with the 3.9 m Anglo-Australian Telescope (AAT) employing the AAOmega multi-object spectrograph (Sharp et al. 2006) and the 2dF fibre feed. The spectra were taken using a 2 arcsec diameter fibre, covering a range between 3 740Å to 8 850Å, with a resolution of 3.2Å. The survey covers a large area of the sky, approximately 286 square degrees, and includes over 340 000 galaxies. This study focuses on the GAMA Phase 1, which collects 192 545 galaxies in 48 square degrees, and a limiting Petrosian magnitude of m19.8 mag in one field and m19.4 mag in the other two (for further details see Baldry et al. (2010), Driver et al. (2011), Hopkins et al. (2013), and Liske et al. (2015)).
On the other hand, the emission lines from GAMA were measured using the Gas AND Absorption Line Fitting (GANDALF) algorithm (Sarzi et al. 2006), which is a simultaneous emission and absorption line fitting algorithm. GANDALF measures and corrects for stellar absorption of the emission lines. All these line measurements include the emission line flux, the equivalent width (EW), and the signal-to-noise ratio (S/N) for the most prominent emission lines, among other parameters (for more details, see Gordon et al. (2018)).
The full GAMA catalogue is matched with two other catalogues, one for stellar masses (Taylor et al. 2011), and one for filaments (Alpaslan et al. 2014a), both from the GAMA survey. The Taylor et al. (2011) catalogue estimates the stellar mass-to-light ratio (/L) from optical photometry using stellar population synthesis based on the Bruzual & Charlot (2003) model. The Alpaslan et al. (2014a) catalogue uses an adapted minimal spanning tree protocol called Friend of Friend (FoF) algorithm (Merson et al. 2013), which groups galaxies based on their proximity to one another, thus creating a network of connected galaxies. It identifies galaxies within a specified linking length and classifies them based on their connection to nearby galaxies.
@@
S/N 3 for H and [O iii] 5007, and S/N 2 for H and [N ii] 6583.
0.3 to ensure the H emission line remains within the optical range, as it shifts out at higher redshifts.
@@
192 545 | ||
18 247 | ||
17 283 |
Therefore, the initial number of galaxies available in the GAMA catalogue for this study was reduced from 192 545 to 17 283 after the filtering process.
Next, we corrected all the emission lines for dust extinction. The extinction curve used in this work is from Cardelli et al. (1989), we assume a case B of recombination (Osterbrock 1989) and = 2.86. Therefore, the corrected flux is given by , where and correspond to corrected and observed flux, respectively. The parameter is the extinction for a given wavelength. When is lower than the theoretical value, then the adopted correction is set to zero.
@@ 2.2 BPT diagram
The Baldwin-Phillips-Terlevich (BPT) diagram (Baldwin et al. 1981; Veilleux & Osterbrock 1987) is commonly used to classify and understand the ionisation sources of emission-line galaxies. This diagram allows us to classify the galaxies from the GAMA survey into three groups: star-forming (SF), active galactic nucleus (AGN), and composite.
The BPT diagram in Figure 1 is plotted using the emission lines of [O iii] 5007, [N ii] 6583, H and H from the 17 283 galaxies selected. We used the curve of Kauffmann et al. (2003) to separate composite from AGN galaxies, while the curve of Kewley et al. (2001) separates composite from SF galaxies.
The estimation of the error bars, shown in Figure 1, was done using Monte Carlo propagation. This method generates simulated data for each variable from a multivariate t-distribution, using the means and the full covariance matrix, which includes the diagonal elements (variances) and the off-diagonal terms (covariances) derived from the uncertainties of the original variables.
@@
13 689 | ||
2 336 | ||
1 258 |
The data in Table 2 indicate that a total of 79% of the sample (13 689 galaxies) consists of SF galaxies, which is the focus of our analysis hereafter.
@@ 2.3 Metallicity and star formation rate estimation
Gas metallicities are estimated using the extinction-corrected fluxes of the SF galaxies, and the calibration method from Pettini & Pagel (2004).
@@
and is in the range -1 1.9. We adopted the O3N2 calibration as it provides reliable metallicity estimates for large samples of star-forming galaxies (Kewley & Ellison 2008). Additionally, since the emission-line ratios involved are close in wavelength, this method is minimally affected by dust extinction (Lara-López et al. 2013).
@@
where the luminosity of the H line, corrected for dust obscuration, is defined by Hopkins et al. (2003) as
where and are the observed fluxes of the H and H lines, respectively, corrected by stellar absorption by GANDALF. The EWHα denotes the equivalent width of the H line in angstroms. Finally, denotes the absolute Petrosian magnitude in the r-band. With these considerations, the representation of the – and –SFR relations can be established.
It is important to note that the fibre size of the AAT could introduce some bias in the estimation of and SFR parameters. This issue is particularly relevant for SFR measurements due to the galaxy’s angular size. If a galaxy is large, its outer regions might not fit entirely within the fibre, potentially leading to incomplete SFR measurements. However, this potential bias is mitigated by the calibration method used by Hopkins et al. (2003), which accounts for such discrepancies in Equation 4. In contrast, the parameter is less significantly affected. According to Kewley et al. (2005), as long as at least 20% of the light of the galaxy falls within the optical fibre, the gas metallicity provides a reasonable approximation of the total value.